OpenStudy (anonymous):

What is the relationship between return and liquidity? * a. Investments with high liquidity have high returns. b. Investments with low liquidity have low returns. c. Investments with high liquidity have low returns. d. There is no relationship between returns and liquidity

OpenStudy (anonymous):

@jabez177

OpenStudy (anonymous):

@Mr.Z @misty1212

OpenStudy (mr.z):

"L=ßW I=Investment or price W= We win this amount L=We lose this amount µ=Probability of Wining 1-µ=Probability of Losing L=Liquidity ß=Liquidity per unit of W When we risk some capital, we find that we´re likely to win money or to lose money. If we buy a lottery ticket, the price we´re willing to pay will depend on the prize obtained. Suppose there are 1000 ticket sold at one dollar each. We buy one ticket. The probability to win will be 1/1000=0.001 one out of one thousand. The prize paid is 100 dollars. Shall we assume the risk ? We´ll only risk our money buying that lottery if P(x)=100 x 0.001 + 0 x 0.999=10 cents If the price is over 10 cents most people won´t buy that lottery. If our investment is over some calculated amount, we shall not risk our money. Of course some people will buy lottery despite that, they´re called risk lovers and some people will not buy the lottery despite a lower price than 10 cents.They are called risk averse. If we buy the lottery for 10 cents we´ll be risk neutral." - azkazk2005 https://answers.yahoo.com/activity/questions?show=PPY6PBEZL3V2IZG4E2XFG7N7X4&t=g

OpenStudy (anonymous):

would the awnser be c then? @Mr.Z

OpenStudy (mr.z):

Yes.